OpenStudy (anonymous):

Algebra 2 question in the thread.

OpenStudy (anonymous):

\[\sqrt[3]{x^2} + (2x+4)/3\sqrt[3]{x} = 5x+4/(3\sqrt[3]{x})\]

OpenStudy (anonymous):

Can someone explain how this works?

OpenStudy (ksaimouli):